OpenStudy (anonymous):

What is the slope of the graph of 14x + 7y = 9? 2 ½ ½ 2

OpenStudy (anonymous):

@claireluna

OpenStudy (welshfella):

write it in the form y = mx + b the value of m will be the slope

OpenStudy (anonymous):

I thought it was 2 but I guess I was wrong.

OpenStudy (anonymous):

I'm sorry, I really can't help with slope.

OpenStudy (welshfella):

7y = -14x + 9 now divide both sides by 7

OpenStudy (anonymous):

9/7=1.2857142857142857142857142857143

OpenStudy (welshfella):

7y / 7 = y now divide -14x and + 9 by 7

OpenStudy (anonymous):

So -14/7= -2

OpenStudy (welshfella):

you need not worry about 9/7 - the slope is the number which goes before x

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

So the answer is -2?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):